OpenStudy (emilyjones284):

Find the work done by the force F=10i+15j in moving an object from P(0,0) to Q(2,3).

OpenStudy (whpalmer4):

What is the distance between P(0,0) and Q(2,3)?

OpenStudy (emilyjones284):

Yeah I think so...(2,3)?

OpenStudy (whpalmer4):

Sorry, \[W = F*d\]

OpenStudy (whpalmer4):

My son is playing some silly video game in the background and it is destroying my brain :-)

OpenStudy (whpalmer4):

Distance between two points \((x_1,y_1)\) and \((x_2,y_2)\) is given by \[d = \sqrt{(x_1-x_2)^2 + (y_1-y_2)^2}\]

OpenStudy (emilyjones284):

Haha no problem (:

OpenStudy (whpalmer4):

easy, because one of the points is (0,0)!

OpenStudy (emilyjones284):

Yeah that wasnt so bad... I got square root of 1

OpenStudy (whpalmer4):

;-( how?

OpenStudy (emilyjones284):

wait was that formula supposed to be x1-y1 and x2-y2?

OpenStudy (whpalmer4):

No, I wrote it like I meant it. It's the Pythagorean theorem, in disguise

OpenStudy (emilyjones284):

Okay so I did square root of (0-0)^2 + (2-3)^2 is that right?

OpenStudy (whpalmer4):

Nope. \((x_1,y_1) = (0,0)\) and \((x_2,y_2) = (2,3)\) Plug into formula\[d = \sqrt{(x_1-x_2)^2 + (y_1-y_2)^2}\]

OpenStudy (whpalmer4):

\[d = \sqrt{(0-2)^2 + (0-3)^2} = \]

OpenStudy (emilyjones284):

Alright now I got square root 13 haha

OpenStudy (emilyjones284):

Woah sorry about that huge space hahaha my computers crazy

OpenStudy (whpalmer4):



OpenStudy (whpalmer4):

just Pythagorean theorem, as I said

OpenStudy (whpalmer4):

\[d=\sqrt{13}\] is correct

OpenStudy (whpalmer4):

Now I think we find the magnitude of the force vector, and multiply it by \(d\) to get our answer

OpenStudy (emilyjones284):

Okay Ill do that now

OpenStudy (emilyjones284):

Alright I got 18

OpenStudy (emilyjones284):

and then when I multiply them together would it just be 18 square root 13

OpenStudy (whpalmer4):

How did you get 18?

OpenStudy (emilyjones284):

square root of 10^2 +15^2?

OpenStudy (emilyjones284):

oh or would it just be square root of 18?

OpenStudy (whpalmer4):

\[\sqrt{10^2+15^2} = \sqrt{100+225} = \sqrt{325} = \sqrt{25*13} = 5\sqrt{13}\]

OpenStudy (whpalmer4):

Now we multiply \(d = \sqrt{13}\) by \(||F|| = 5\sqrt{13}\) and get \(W = \) drum roll please....a hush falls over the crowd...will she get it?

OpenStudy (emilyjones284):

Umm 65? :D

OpenStudy (whpalmer4):

the crowd goes nuts!!!

OpenStudy (emilyjones284):

Yayyyyy thank you! I think my brain is fried but I understand everything a lot better now! (:

OpenStudy (whpalmer4):

Come on back, let's finish off the cube root problem...

OpenStudy (emilyjones284):

Okay haha Ill try to do it quick because I have to go to bed :P